### v1

# Binary-Report Peer Prediction for Real-Valued Signal Spaces

Abstract
Theoretical guarantees about peer prediction mechanisms typically rely on the discreteness of the signal and report space. However, we posit that a discrete signal model is not realistic: in practice, agents observe richer information and map their signals to a discrete report. In this paper, we formalize a model with real-valued signals and binary reports. We study a natural class of symmetric strategies where agents map their information to a binary value according to a single real-valued threshold. We characterize equilibria for several well-known peer prediction mechanisms which are known to be truthful under the binary report model. In general, even when every threshold would correspond to a truthful equilibrium in the binary signal model, only certain thresholds remain equilibria in our model. Furthermore, by studying the dynamics of this threshold, we find that some of these equilibria are unstable. These results suggest important limitations for the deployment of existing peer prediction mechanisms in practice.
1 Introduction
Eliciting high-quality information from agents without verification is a well-studied problem, motivated by settings such as peer grading or data labeling where the ground truth is unknown, subjective, or hard to acquire. Peer prediction mechanisms aim to incentivize truthful behavior of agents by paying them based on how their reports correlate with other submissions, without the need for ground truth. Agents are typically modeled as receiving a signal, for example $H$ (“high”) or $L$ (“low”) for the quality of an essay, from a fixed joint probability distribution $P$ unknown to the mechanism. Upon seeing their signal, agents update their beliefs and, under the right conditions, are incentivized to truthfully report the signal they receive.
In practice, however, the signal space of information that agents receive is much richer than the small number of categories typically studied in the literature. An essay contains far more information than a simple $H$ or $L$ impression. As a result, one would expect the same agent to form a range of posterior beliefs based on these more nuanced impressions. What are the implications of this more nuanced signal space? [^1] One implication which we do not focus on is the well-known issue of spurious correlation, such as adopting a strategy of conditioning one’s report on the first word of the essay; see § 7. Existing results often rely on a clear separation in the joint distribution $P$ to control the incentives, making it unclear how one would expect these mechanisms to perform in practice.
To address this question, we introduce and study a model of binary-report peer prediction for real-valued signal spaces. In this model, agents receive a real number as their signal and then based on it select a report of $H$ or $L$. We focus on the natural class of threshold strategies, where an agent reports $H$ if and only if their signal exceeds a particular threshold. For example, an agent may report $H$ if they deem the quality of an essay to be above 0.7, and $L$ otherwise.
We generally find that mechanisms which are truthful under the original binary signal model fail to yield correct incentives under the more nuanced model with real-valued signals. Intuitively, agents with signals near the threshold may have an incentive to misreport, as the separation that enforces truthfulness breaks down. We give necessary and sufficient conditions for a threshold to be an equilibrium, i.e., not suffer from this issue, in a variety of peer prediction mechanisms. Furthermore we study dynamics arising if a small fraction of agents with the greatest incentive to do so change their report, causing the threshold to slowly shift. We then distinguish stable equilibria under our dynamics as reasonable to arise naturally, from unstable equilibria.
1.1 Motivating example: Output Agreement (OA)
Consider Output Agreement (OA), which provides a reward of 1 if an agent’s report agrees with a peer and 0 otherwise [Von Ahn and Dabbish, 2004, 2008]. In our more nuanced, real-valued signal model, each agent $i$ recieves a signal we model as a random variable $X_{i}=Z+Z_{i}$ where $Z$ and $Z_{i}$ are both unit normals with $Z$ shared by all agents and $Z_{i}$ unique to agent $i$. In a peer grading setting the signals $X_{i}$ might capture a fine-grained assessment of the quality of an assignment.
In contrast, the traditional signal structure for OA is usually taken to be finite, so take the simplest binary case where the set of signals is $\{H,L\}$. The mechanism designer has some intended mapping between the real-valued and binary signal spaces. Given the interpretation of $H$ and $L$ as “high” and “low”, it is natural for this mapping to be determined by a real-valued threshold $\tau$, with all values above the threshold corresponding to $H$ and all below it corresponding to $L$. In peer grading this naturally corresponds to an intution that a “good-enough” assignment might be graded as satisfactory while a poor one would be deemed unsatisfactory.
Let us suppose first that the mechanism designer announces a desired threshold of $\tau=0$. In the binary signal model with this threshold and distribution (i.e. agents only receive a signal of $H$ or $L$ as determined by $\tau$), OA is truthful because agent $i$ receiving $H$ makes it more likely that all other agents received $H$ than $L$ and vice versa ($\Pr[H|H],\Pr[L|L]>0.5$).
In the real-valued case, we can no longer speak of OA as truthful because it is no longer a direct revelation mechanism, but we can ask whether reporting as intended by $\tau$ is a (Bayes-Nash) equilibrium. It turns out $\tau=0$ is such an equilibrium. For example, if $x_{i}=0.11$, we have $\Pr[X_{j}>0|x_{i}]\approx 0.52>0.5$, so reporting $H$ as intended is indeed the best response.
Now suppose the desired threshold were $\tau=0.1$. In the binary signal model, $\Pr[H|H]=\Pr[X_{j}>0.1|X_{i}>0.1]\approx 0.64$ and $\Pr[L|L]\approx 0.69$ are both greater than $0.5$, so OA remains truthful. Now in the real-valued model consider an agent receiving a signal of $x_{i}=0.11$. As $x_{i}$ is just above the threshold, the intended report is $H$. But $\Pr[X_{j}>0.1|x_{i}]\approx 0.48<0.5$. Thus the best response is $L$, so reporting according to $\tau$ is not an equilibrium. The above analysis will give the same qualitative result for any finite nonzero $\tau$, intuitively because with a positive threshold any agent receiving a signal slightly above the threshold will think that the average agent will have a signal below it and vice versa for negative thresholds. Thus, the only finite initial threshold $\tau$ for which reporting according to $\tau$ is an equilibrium is $\tau=0$.
Given that $\tau=0.1$ is not an equilibrium, what should we expect to happen if the mechanism designer announces it? We saw that agents with signals $x_{i}$ slightly above $0.1$ have an incentive to report $L$ instead of $H$. This incentive weakens as $x_{i}$ grows, so it is natural to expect only a small fraction of agents to deviate from the intended strategy, specifically those with signals closest to but above $0.1$. But effectively, this raises our threshold from $\tau_{0}=0.1$ to some higher value, say $\tau_{1}=0.13$. Since the same argument applies to our new threshold, in future tasks further deviations could lead to $\tau_{2}=0.16$. Thus, we should expect that, over time, the system will converge toward the “trival” equilibrium at $\tau=\infty$ where agents report $L$ regardless of their signal. The same logic shows that announcing $\tau<0$ will lead toward the trivial “always $H$” equilibrium at $\tau=-\infty$.
From a dynamical systems perspective, the above shows that the $\tau=0$ equilibrium is unstable and can only be realized by starting exactly at it. Since there are no other finite equilibria, any variation or small error in the model will inexorably lead to a threshold that tends toward infinity. Given the inherent noisiness of many peer prediction applications, we argue similar to Shnayder et al. [2016] that the only reasonable equilibra in practice are stable ones. Stability is thus a crucial equilibrium refinement for peer prediction.
To summarize, the standard analysis of OA in the binary signal model concludes that the mechanism has desirable incentive properties as long as agents believe others are more likely to share their signal than have the opposite. In contrast, by introducing our more nuanced real-valued signal model, we have seen that the only reasonable outcome to expect from OA is a “trivial” equilibrium where no information is gained. Thus, at least in this example, this shift in perspective takes OA from useful to entirely useless.
1.2 Results
[…]
In the binary signal setting, a multi-task mechanism proposed by Dasgupta and Ghosh [2013] (DG) fixes the uninformative equilibria problem in OA. The DG mechanism adds a penalty corresponding to the empirical frequency of agents’ reports across other tasks, effectively discouraging blind agreement and leading to strictly worse payments in uninformative equilibria. Under our richer signal model, we provide necessary and sufficient conditions for a threshold to be an equilibrium. Under some conditions like monotonicity, these equilibria are stable, while the uniformative equilibria are unstable. We conclude that DG is potentially useful but may only allow limited control over the equilibrium threshold, i.e., what $H$ and $L$ end up meaning.
We perform similar analyses for two other more complex mechanisms, Determininant Mutual Information (DMI) [Kong, 2020, 2024] and the Robust Bayesian Truth Serum (RBTS) [Witkowski and Parkes, 2012]. Under DMI, an interesting consequence of our model is the benefit of coordinating reports across tasks. For all four mechanisms we study we augment our general analysis with a worked example of the Gaussian case and numerical illustrations of more complex distributions that exhibit skewness and multimodality.
In general, our results show that realistic behavior under common peer prediction mechanisms departs significantly from the traditional discrete-signal theory. With a richer signal space, agents will essentially redefine the meaning of the binary $H$ and $L$ signals based on community norms. This paper therefore serves as a launching point for uncovering how behavior changes under peer prediction in the real world. In particular, our results will help practitioners better understand the implications of their choice of mechanism.
1.3 Related Work
For a general overview of peer prediction, we direct the reader to Faltings [2023], Lehmann [2024], Frongillo and Waggoner [2024]. Most relevant to our work is the distinction between peer prediction mechanisms which are minimal, in that agents only report their signal (as in OA), and those where additional information about agent beliefs is reported (e.g. the agent’s posterior belief about the report of another agent as in the Bayesian Truth Serum [Prelec, 2004] and its variants). Another important dimension is whether the mechanism is for a single task or a larger mult-task collection where reports from unrelated tasks can be compared to improve incentives. We examine two multi-task mechanisms (DG and DMI) and a non-minimal one (RBTS). While more robust than OA, our results show that neither of these structures avoids the issues our work raises.
[…]
Looking closer, two key differences between our work and Shnayder et al. [2016] are the signal model and strategic model. For the signal model, they consider a discrete signal space matching the report space, as usual. For the strategic model, they consider agents learning whether to switch to another strategy in a discrete set, whereas we allow agents to choose any threshold strategy. As a result, our results differ substantially in some cases: for OA, the truthful equilibrium is unstable even in a technical sense (measure zero basin of attraction), and for DG, “truthfulness” as given by the mechanism designer’s desired threshold need not be an equilibrium.
We note that the model and tools in our paper also bear some resemblance to past work in the epistemic democracy literature. Specifically, Duggan and Martinelli [2001] and Meirowitz [2002] study a common-value voting setting, where there is a ground truth “correct” alternative that all agents prefer, and they each receive signals about which alternative that is. As in our paper, the alternative space is binary while the signal space is continuous. The authors similarly identify threshold equilibria, where agents map their signals to a vote for either alternative according to a cutoff point, and prove that under some conditions on the continuity and monotonicity of the prior densities over signals in each world state, there is a unique symmetric Bayes-Nash equilibrium characterized by a threshold. We note that though the tools are similar (equilibria are characterized by evaluating conditional probabilities at the threshold), the setting and goals of the papers are quite different. (The authors aim to show that the “right” alternative is chosen with high probability over the limit of the number of agents.)
2 Output Agreement
We begin by analyzing the popular peer prediction mechanism Output Agreement (OA). Output Agreement is a minimal, single task mechanism: each agent $i$ submits a report for a task and receives a positive payment if their report matches that of another agent $j$ on the same task. In the binary signal model, truthful reporting forms a correlated equilibrium under some assumptions about the information structure. Specifically, truthfulness holds under strong diagonalization, where the conditional probability of another signal is maximized by the signal being conditioned on ($\Pr[H\mid H]>\Pr[L\mid H]$ and $\Pr[L\mid L]>\Pr[H\mid L]$). However, uninformative equilibria where agents misreport their information also exist. Specifically, agents can coordinate to all submit the same report and each receive a maximum payment. Nonetheless, Output Agreement remains popular because of its simplicity. We thus aim to characterize OA under our richer signal space model to help practitioners understand how behavior occurs in the real world.
2.1 Model
We begin by introducing the model that, with minor variations to accommodate the form of different peer prediction mechanisms, will be used througout the paper. There are $n$ agents; each agent $i$ receives a signal $X_{i}\in\mathbb{R}$ representing information gained about a shared task. The signals are drawn from a joint distribution $\mathcal{D}$ that is symmetric in the sense that each agent $i$ has (1) the same marginal $F(x)=\Pr[X_{i}=x]$, and (2) the same posterior distribution $\beta(x)=\Pr[X^{\prime}=\cdot\mid X_{i}=x]$. Identical marginals are realistic when the signals are exchangeable, e.g. when they have the same conditional distribution over some latent variable $\theta$.
Let the report space be =$\mathcal{R}=\{L,H\}$. Each agent $i$ calculates a report $r_{i}\in\mathcal{R}$ for the task according to a deterministic strategy $\sigma_{i}:\mathbb{R}\to\mathcal{R}$ mapping from signal to report space. The Output Agreement mechanism pays each agent $i$ an amount $M_{\textrm{OA}}(r_{i},r_{j})$ as a function of reports $r_{i},r_{j}$, where
$M_{\textrm{OA}}(r_{i},r_{j})=\mathbf{1}[r_{i}=r_{j}].$
Thus the (ex-interim) expected utility for playing strategy $\sigma_{i}$ when $j$ plays according to $\sigma_{j}$ is
$U_{i}(\sigma_{i},\sigma_{j},x)=\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}\mathbf{1}[\sigma_{i}(x)=\sigma_{j}(x^{\prime})].$ (1)
We consider a natural class of strategies, threshold strategies, where agent $i$ reports $H$ if and only if their signal $x$ satisfies $x\geq\tau$ for some fixed threshold $\tau\in\mathbb{R}\cup\{\pm\infty\}$. That is, for some $\tau$,
$\sigma^{\tau}(x)=\begin{cases}L,&x\leq\tau\\ H,&x>\tau.\end{cases}$
We will often denote a strategy directly by its threshold $\tau$ when it is clear. Thresholding is a natural way to assign semantic meaning to the discrete labels $H$ and $L$ according to the continuous signal space. Moreover, in many settings the mechanism designer themselves may announce a threshold to establish norms that they would like agents to follow. For example, in peer grading a teacher may establish a higher threshold to encourage a high bar for good marks.
2.2 Equilibrium Characterization
We are interested in characterizing threshold strategies which are symmetric Bayes-Nash equilibria. That is, equilibria where each agent $i$ commits to the same threshold strategy $\sigma^{\tau}$. For brevity, we refer to these as threshold equilibria. Symmetric strategies are natural in our model since agents share the same ex-ante expected utilities.
Definition 1.
A threshold strategy $\sigma^{\tau}:\mathbb{R}\to\mathcal{R}$ is a threshold equilibrium under OA if
$\forall x\in\mathbb{R},\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}\mathbf{1}[\sigma^{\tau}(x)=\sigma^{\tau}(x^{\prime})]\geq\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}\mathbf{1}[\overline{\sigma^{\tau}(x)}=\sigma^{\tau}(x^{\prime})],$ (2)
where $\overline{H}=L,\overline{L}=H.$
In the case where the signal $x$ the agent receives satisfies $x\leq\tau$. Condition (2) simplifies to
$\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}\mathbf{1}[x^{\prime}\leq\tau]\geq\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}\mathbf{1}[x^{\prime}>\tau],\mbox{or equivalently }\Pr[X^{\prime}\leq\tau\mid X=x]\geq 1/2.$
Similarly if $x>\tau$, (2) simplifies to $\Pr[X^{\prime}\leq\tau\mid X=x]\leq 1/2$. Let $P(\tau;x)=\Pr[X^{\prime}\leq\tau\mid X=x]$ be the probability another agent reports $L$ conditioned on seeing a signal $x$, for a fixed threshold $\tau$. As a summary, then, Condition (2) is equivalent to
$\forall x\leq\tau,P(\tau;x)$ $\geq 1/2,$ (3)
$\forall x>\tau,P(\tau;x)$ $\leq 1/2.$ (4)
Results.
Under these conditions, we first show that the thresholds $\tau^{*}=\pm\infty$ are always equilibria under Output Agreement.
Proposition 1.
$\tau^{*}=\pm\infty$ are both always threshold equilibria under OA.
Proof.
Let $\tau^{*}=\infty$. Then for all signals $x$ such that $x\leq\tau^{*}$, $P(\tau^{*};x)=1\geq 1/2$ (since any $X^{\prime}\in\mathbb{R}$ satisfies $X^{\prime}\leq\tau^{*}$ with probability one). Meanwhile, Statement 4 is vacuous since no signal $x\in\mathbb{R}$ satisfies $x>\infty$. A similar argument follows for $\tau^{*}=-\infty$: for all signals $x>\tau^{*}$, $P(\tau^{*};x)=0\leq 1/2$, while Statement 3 is vacuous since no signal $x\in\mathbb{R}$ satisfies $x<-\infty$. ∎
These thresholds exactly correspond to uninformative equilibria in the binary setting: if $\tau^{*}=\infty$, agents always report $L$, and if $\tau^{*}=-\infty$, agents always report $H$. Under some smoothness conditions on the predictive posterior function, we can furthermore characterize all non-infinite threshold equilibria. To do so, we define the function $G(x)=P(x;x)=\Pr[X^{\prime}\leq x\mid X=x]$ as the probability of another signal lying below $x$, conditioned on $x$. In Output Agreement, we find that threshold equilibria are characterized by thresholds $\tau$ such that $G(\tau)$ crosses 1/2. More precisely, we give the following necessary and sufficient conditions.
Theorem 1.
Let finite threshold $\tau$ be given and $P(\tau;x)$ be continuous in $x$. If $\tau$ is a threshold equilibrium under the OA mechanism then $G(\tau)=1/2$. Conversely, if $G(\tau)=1/2$ and either (a) $P(\tau;x)$ is monotone decreasing in x or (b) $P(\tau;x)-1/2$ has a single crossing of 00 from positive to negative then $\tau$ is a threshold equilibrium under the OA mechanism.
Proof.
For necessity, assume that $\tau$ is a threshold equilibrium, so Equations (3) and (4) hold. Since $P(\tau;x)$ is continuous over $x$, we must have $\lim_{x\to\tau^{+}}P(\tau;x)=\lim_{x\to\tau^{-}}P(\tau;x)=1/2.$ But $\lim_{x\to\tau^{+}}P(\tau;x)\leq 1/2$ and $\lim_{x\to\tau^{+}}P(\tau;x)\geq 1/2$, so we must have $P(\tau;\tau)=G(\tau)=1/2$.
For sufficiency, assume $G(\tau)=1/2$. Equivalently $P(\tau;\tau)-1/2=0$, so (a) implies (b) and the single crossing is at $\tau$. Thus assume (b) holds and take some $x>\tau$. By single crossing, $P(\tau;x)\leq P(\tau;\tau)=G(\tau)=1/2$. Similarly, if $x\leq\tau$, $P(\tau;x)\geq P(\tau;\tau)=1/2$. This establishes Equations (3) and (4), so $\tau$ is a threshold equilibrum. ∎
For intuition for the conditions, observe that the derivative of expected utility with respect to agent $i$’s choice of threshold $\tau$ in Equation (1) is $f(x)(2G(x)-1)$, where $f$ is the p.d.f. of $F$ and thus non-negative. Thus the necessary condition corresponds to the first order condition of optimizing the best response threshold and the sufficient conditions ensure this optimization is concave and quasiconcave respectively. However, the proof provided establishes optimality in the space of all possible strategies (which is required to be a threshold equilibrium), not merely that the chosen threshold is the optimal threshold.
More broadly, this result shows that “truthfulness” in the sense of following the strategy prescribed by the mechanism designer is fragile in this setting. Only a sharply limited set of thresholds can be implemented in equilibrium—in some cases a single one. This means the mechanism designer’s ability to choose the “meaning” of $H$ and $L$ can be quite limited in practice.
2.3 Dynamics
We have characterizations of both the uninformative infinite equilibria and more interesting finite equilibria. Which are more likely to occur? To answer this question, we turn to modeling the dynamics of the system. We are interested in studying how agents will behave under our model over time, as a tool for equilibrium refinement. That is, we want to know what threshold agents will land on as they continue to grade new batches of papers or submit new labels. Since it is likely that the mechanism shares empirical results before all agents equilibriate, we consider a dynamic setup where a small fraction of agents are able to best respond at each time step.
Formally, consider the following discrete best response dynamic: the mechanism publishes an initial threshold $\tau(0)$, the “norms” with which agents should interpret their signals as $H$ and $L$. At each time step, a small, random fraction $\Delta t$ of agents are able to commit to a new best response threshold strategy $\hat{\tau}(t)$ against the current threshold $\tau(t)$. Meanwhile, the rest of the agents stick to the threshold they used previously, whether that be a previous best response or the initial $\tau(0)$. Note that since $\hat{\tau}(t)$ is a best response, its calculation depends on (1) the peer prediction mechanism’s payment scheme and (2) the current threshold $\tau(t)$. Finally, the empirical frequency of reports is updated to reflect this new mixture of thresholds, and the process repeats. This dynamic corresponds to the following equation:
$\tau(\Delta t)$ $=(1-\Delta t)\tau(0)+\Delta t\hat{\tau}(0)$
$\tau(2\Delta t)$ $=(1-\Delta t)\tau(\Delta t)+\Delta t\hat{\tau}(\Delta t)$
$\ldots$
$\tau(t+\Delta t)$ $=(\Delta t)\hat{\tau}(t)+(1-\Delta t)\tau(t).$ (5)
We note that $\hat{\tau}(k\Delta t)$ is technically a best response to the mixture of thresholds at the previous time step $k-1$, rather than the probability distribution over thresholds. Ignoring this for the moment and taking the limit as $\Delta t\to 0$ of the discrete system in Equation 5, we end up with the following continuous best response dynamic:
$\dot{\tau}=\hat{\tau}-\tau.$ (6)
While not, strictly speaking, the correct update from best response dynamics, this update rule has the appealing form of the threshold slowly drifting toward the best response. Alternatively, since agents near the threshold have the strongest incentive to change their strategy, this update rule could be intepreted as a form of quantal response. In any case, our focus in this work is on what sort of equilibria we should expect to end up at rather than the exact process the system takes to get there, so this choice of dynamics seems simple, natural, and amenable to analysis.
We begin our study of OA under the dynamics described by Equation (6) with the following observation. A nice property of Output Agreement (which relates to our sufficient conditions from Theorem 1) is that for a fixed threshold $\tau$, under decreasing monotonicity of $P(\tau;x)$ the best response over all strategies is always a threshold even out of equilibrium.
Proposition 2.
Assume for a fixed threshold $\tau$ that $P(\tau;x)$ is strictly decreasing and continuous over $x$. If all agents are playing according threshold strategy $\tau$, the unique best response of an agent across all strategies $\sigma:\mathbb{R}\to\mathcal{R}$ is to play according to threshold strategy $\hat{\tau}$ satisfying $P(\tau;\hat{\tau})=1/2$.
Proof.
If $\tau$ is the current threshold strategy that all other agents are following, an agent who receives signal $x$ will best respond with $H$ if $\Pr[X\leq\tau\mid X=x]\leq 1/2$, and $L$ otherwise. Since $P(\tau;x)$ is strictly decreasing and continuous over $x$, there will be a unique point $\hat{\tau}$ such that $P(\tau;\hat{\tau})=1/2$, with $P(\tau;x)\geq 1/2$ for $x\leq\hat{\tau}$ and $P(\tau;x)\leq 1/2$ for $x>\hat{\tau}$. This threshold $\hat{\tau}$ thus corresponds to the best response strategy. ∎
We can now use concepts from dynamical systems and characterize equilibria as stable or unstable according to properties of the function $G$. This distinction captures the behavior of the dynamics near the equilibrium: do they drive the system toward the equilibrium (stable) or away from it (unstable)? More precisely, an equilibrium is (locally) stable if a sufficiently small perturbation yields dynamics that always return to it while it is unstable if all perturbations diverge from it.
We argue that only stable equilibria are reasonable to find in practice. Otherwise any small error in the choice of $\tau^{*}$ will fail to yield the desired equilibrium as the dynamics push away from it. Since the equilibria depend on the joint distribution $\mathcal{D}$, which will typically not be perfectly known to the mechanism designer, such small errors are to be expected. As a result, we view stability as an important equilibrium refinement in the context of peer prediction.
Theorem 2.
Assume for all $\tau\in\mathbb{R}$ that $P(\tau;x)$ is strictly decreasing and continuous over $x$. Then if $G(\tau)$ is strictly increasing at equilibrium point $\tau^{*}$, $\tau^{*}$ is unstable, while if it is instead strictly decreasing $\tau^{*}$ is stable.
Proof.
Consider the dynamics at time step $t$, with the current threshold $\tau(t)$. Since we are considering a fixed time $t$, we refer to $\tau(t)$ as $\tau$ throughout the proof.
We consider what happens when an agent receives a signal exactly at $\tau$. There are three cases to consider. In the first, $G(\tau)=1/2$; then the expected utility of reporting $L$ is $\Pr[X^{\prime}\leq\tau\mid X=\tau]=1/2$, while the expected utility of reporting $H$ is $\Pr[X^{\prime}>\tau\mid X=\tau]=1/2$; thus the agent is indifferent between reporting $L$ or $H$. Moreover, the best response $\hat{\tau}$ is exactly $\tau$. It follows the system is at an equilibrium of the dynamics since $\dot{\tau}=\hat{\tau}-\tau=0$.
In the second case, $G(\tau)>1/2$. Then reporting $L$ is strictly preferred to reporting $H$. Since $P(\tau;\hat{\tau})=1/2$, by decreasing monotonicity we have $\hat{\tau}>\tau$. Thus $\dot{\tau}=\hat{\tau}-\tau>0$. In the third case, $G(\tau)<1/2$. Then reporting $H$ is strictly preferred to reporting $L$. Since $P(\tau;\hat{\tau})=1/2$, by decreasing monotonicity we have $\tau>\hat{\tau}$. Thus $\dot{\tau}=\hat{\tau}-\tau<0$.
If $G(\tau)$ is strictly increasing at $\tau^{*}$, then at any perturbed point $\tau$ to the left of $\tau^{*}$ we have $G(\tau)<1/2$ and $\dot{\tau}<0$; and at any perturbed point $\tau$ to the right of $\tau^{*}$ we have $G(\tau)>1/2$ and $\dot{\tau}>0$. It follows that $\tau^{*}$ is unstable. The same logic implies stability in the strictly decreasing case. ∎
In most reasonable settings, one would expect $G(x)$ to limit toward 1 as $x$ approaches $\infty$, and toward 0 as $x$ approaches $\infty$- ∞. Specifically, as a signal $x$ becomes increasingly small, the probability another agent receives a signal smaller than $x$ should approach 0. Meanwhile, as a signal $x$ grows large, the probability another agent receives a signal smaller than $x$ should approach 1. As the quality of essay increases in a normally distributed class, for example, a grader would believe that the mass of essays below that quality should increase proportionally.
We show formally in § A that under such limit behavior of $G$, uninformative equilibria at $\tau^{*}=\pm\infty$ are stable in OA, while at least one non-infinite equilibrium exists which is unstable. In fact, one would expect in smooth, unimodal settings, as in § 2.4, that $G$ is monotone increasing. In this case, there are three equilibria, and by Theorem 2, the internal equilibrium is unstable while the uninformative equilibria are stable.
So what does this instability of internal equilibria mean for OA? When signals correspond monotonically to quality of an essay or task, agents will naturally move toward uninformative equilibria. These dynamics only further the fragility of truthfulness in OA: unless the threshold $\tau$ exactly balances out the conditional probabilities that other agents will report $H$ or $L$, agents will inevitably and increasingly misreport until they reach an uninformative consensus. There are multimodal Bayesian examples where several internal equilibria exist and some are thus topologically required to be locally stable (see § 6). However, as long as $G$ remains monotone increasing at extreme signals, we expect uninformative equilibria to be stable by our analysis in § A.
Figure 1: Plots of $F$, $G$, and $Q$ in the Gaussian model (§ 2.4) for the parameters $a=b=1$.
Figure 2: Plot of $G-F$ in the Gaussian model (§ 2.4), for parameters $a=b=1$. $G(x)-F(x)$ is decreasing at $x=0$, meaning the corresponding equilibrium in DG is stable under the dynamics.
2.4 A Gaussian Model
Suppose each agent $i$ receives a noisy signal $X_{i}$ from a Gaussian distribution, with the noise also normally distributed. That is, $X_{i}=aZ+bZ_{i}$, where $Z\sim N(0,1)$, $Z_{i}\sim N(0,1)$, and $a,b\in\mathbb{R}_{>0}$. This model e.g. realistically captures peer grading settings where there is a more fine-grained evaluation of essays anchored in a standard bell-curve distribution. It follows that the marginal distribution for each agent is $F(x)=N(x\mid 0,a^{2}+b^{2})$. This model allows for a continuum of agents with the same $F$, so the setting is consistent with our dynamics.
The correlation coefficient for two signals $X,X^{\prime}$ is $\rho=\frac{a^{2}}{a^{2}+b^{2}}$. Fixing any agent, the predictive posterior distribution upon seing $x$ is $\Pr[X^{\prime}=x^{\prime}\mid X=x]=N(x^{\prime}\mid\hat{\mu}(x),\hat{\sigma}^{2})$, where $\hat{\mu}(x)=\rho x$ and $\hat{\sigma}^{2}=b^{2}(1+\rho).$ Denoting the CDF of a standard Normal distribution by $\Phi$, it follows that for a fixed $\tau$,
$P(\tau;x)=\Phi\left(\frac{\tau-\rho x}{b\sqrt{1+\rho}}\right).$ (7)
We can now also write
$F(x)=\Phi\left(\frac{x}{\sqrt{a^{2}+b^{2}}}\right)=\Phi\left(\frac{\sqrt{\rho}}{a}\;x\right)\mbox{ and }G(x)=\Phi\left(\frac{x-\rho x}{b\sqrt{1+\rho}}\right)=\Phi\left(\frac{(1-\rho)}{b\sqrt{1+\rho}}\;x\right)~{}.$ (8)
We then immediately observe existence of equilibria according to our general results:
Corollary 1.
[…]
Proof.
Existence of equilibria at $\tau=\pm\infty$ follows from Proposition 1. Note that $P(\tau;x)$ is strictly monotone decreasing and continuous in $x$ so that Theorem 1 applies. As $G(0)=\Phi(0)=\tfrac{1}{2}$, $\tau=0$ is an equilibrium for OA. Note also that $\Phi$ is strictly monotone, so $G(x)$ only crosses 1/2 once and $\tau=0$ is therefore the only non-infinite equilibrium (see a visualization in Figure 2). ∎
Now, consider the stability of these equilibria.
Corollary 2.
[…]
Proof.
Let $c_{G}(\rho)=\frac{(1-\rho)}{b\sqrt{1+\rho}}$ be the coefficient of $x$ in $G(x)$. Then $G^{\prime}(x)=c_{G}(\rho)\phi(c_{G}x)$ for $\phi$ the PDF of the standard Normal, and so $G^{\prime}(0)=c_{G}(\rho)\phi(0)>0$. By Theorem 2, then, $\tau=0$ is unstable. Since 00 is the only non-infinite equilibrium, the equilibria $\tau=\pm\infty$ are stable by topological necessity. ∎
For any finite starting point $\tau(0)\neq 0$, then, we have $\lim_{t\to\infty}\tau(t)=\pm\infty$. Therefore, we find in settings where graders receive a noisy, normally distributed version of information about an essay, Output Agreement incentivizes agents to stabilize at an equilibrium where they all submit the same report. This uninformative consensus is in contrast to the truthful equilibrium guarantees of the binary signal model. [^2] In the Gaussian model, strong diagonalization holds within a reasonable range of thresholds around 0, depending on $\rho$.
3 Dasgupta-Ghosh
We next consider the mechanism proposed by Dasgupta and Ghosh [2013], which we will refer to as the Dasgupta-Ghosh (DG) mechanism. DG is a multi-task mechanism: agents participate in a set of peer prediction tasks, and their payment is a function of their reports across these tasks. For a given task, agent $i$ receives a payment of one if their report $r_{i}$ matches that of another agent $j$ on that task, as in Output Agreement. However, they are also penalized if their report matches that of the other agent on some unrelated task agent $i$ did not complete, with the intution that agents should be paid only for their performance over and above what would be expected by chance. [^3] There have been a variety of ways considered to calculate this penalty. This version corresponds to their $d=1$ option. In the binary signal setting, this penalty ensures that uninformative equilibria in OA where all agents agree to report H or L now have zero value, thereby ensuring strategic agents are not incentivized toward them.
Formally, if $r_{i}$ agent $i$’s report on the task, $r_{j}$ is the peer report on the task, and $r_{k}$ is the peer report on a distinct task,
$M_{\textrm{DG}}(r_{i},r_{j},r_{k})=\mathbf{1}[r_{i}=r_{j}]-\mathbf{1}[r_{i}=r_{k}].$
Assume the task $r_{k}$ comes from is i.i.d. to the task being scored and, as with our analysis of OA, all agents are using symmetric strategies. Then agent $i$’s expected penalty is $\mathop{\mathbb{E}}\mathbf{1}[r_{i}=\sigma(X)]$. That is, the expected penalty is exactly the prior probability of a signal that leads to a report of $r_{i}$. We denote this as
$\pi_{L}=\Pr[\sigma(X)=L]=\Pr[X\leq\tau]=F(\tau),$
with $\pi_{H}=1-\pi_{L}$. Therefore, we instead work with the following formulation of DG which has the same expected payment:
$M_{\textrm{DG}}(r_{i},r_{j},\pi_{L})=\mathbf{1}[r_{i}=r_{j}]-\pi_{r_{i}}.$
Despite the nature of DG as a multi-task mechanism, this means that (under our assumption of symmetry and if other agents use the same strategy across tasks) we can essentially analyze each task in isolation. More formally, this means we can adopt the same model from § 2.1 we used to analyze OA, changing only the payment rule and updating the quantities that derive from it accordingly.
3.1 Equilibrium Characterization
For DG, the (ex-interim) expected utility for playing strategy $\sigma_{i}$ depends on $\sigma_{j}$ both directly in the bonus for agreement and indirectly through $\pi_{L}$:
$U_{i}(\sigma_{i},\sigma,x)=\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}\mathbf{1}[\sigma_{i}(x)=\sigma(x^{\prime})]-\pi_{\sigma_{i}(x)}.$ (9)
The condition for a symmetric Bayes-Nash equilibrium can then be written as
$\forall x\in\mathbb{R},\;\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}\mathbf{1}[\sigma(x)=\sigma(x^{\prime})]-\pi_{\sigma(x)}\geq\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}\mathbf{1}[\overline{\sigma(x)}=\sigma(x^{\prime})]-\pi_{\overline{\sigma(x)}}.$
Since $\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}\mathbf{1}[\sigma(x)=\sigma(x^{\prime})]=1-\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}\mathbf{1}[\overline{\sigma(x)}=\sigma(x^{\prime})]$, this condition simplifies to
$\forall x\in\mathbb{R},\;\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}\mathbf{1}[\sigma(x)=\sigma(x^{\prime})]\geq\pi_{\sigma(x)}.$ (10)
For a threshold equilibrium, Condition 10 is equivalent to:
$\forall x\leq\tau,P(\tau;x)\geq\pi_{L}$ (11)
$\forall x>\tau,P(\tau;x)\leq\pi_{L}.$ (12)
Equilibrium results.
Now we have the tools to analyze threhsold equilibria in DG; we leave proofs to § B, since they largely follow the same logic as our proofs for OA. We first observe that uninformative strategies $\tau^{*}=\pm\infty$ remain equilibria in the DG mechanism.
Proposition 3.
$\tau^{*}=\pm\infty$ are both always threshold equilibria under DG.
We can also provide necessary and sufficient conditions for a finite threshold equilibrium for DG. Both their form and proof follow the same logic we saw for OA, adapted to account for the additional $\pi_{L}$ term also depending on the choice of threshold.
Theorem 3.
Let finite threshold $\tau$ be given and $P(\tau;x)$ be continuous in $x$. If $\tau$ is a threshold equilibrium under the DG mechanism then $G(\tau)=F(\tau)$. Conversely, if $G(\tau)=F(\tau)$ and either (a) $P(\tau;x)$ is monotone decreasing in x or (b) $P(\tau;x)-\pi_{L}$ has a single crossing of 00 from positive to negative then $\tau$ is a threshold equilibrium under the DG mechanism.
3.2 Dynamics
We consider stability of equilibria for DG under the dynamics described by Equation (6). Again, we leave proofs in this section to § B. Note that as in OA, when everyone else is playing according to $\tau$, there exists a unique best response which is also a threshold strategy.
Proposition 4.
Assume that $P(\tau;x)$ is strictly decreasing and continuous over $x$ for a fixed threshold $\tau$. If all agents are playing according threshold strategy $\tau$, the unique best response of an agent across all strategies $\sigma:\mathbb{R}\to\mathcal{R}$ is to play according to threshold strategy $\hat{\tau}$ satisfying $P(\tau;\hat{\tau})=F(\tau)$.
Now, as in OA, we can characterize the stability of equilibria.
Theorem 4.
Assume for all $\tau\in\mathbb{R}$ that $P(\tau;x)$ is strictly decreasing and continuous over $x$. Then if $G(\tau)-F(\tau)$ is strictly decreasing at equilibrium point $\tau^{*}$, $\tau^{*}$ is stable. Similarly, if it is strictly increasing $\tau^{*}$ is stable.
The difference between $G$ in the OA case and $G-F$ here is important. We saw that $G$ was strictly increasing in natural settings like the Gaussian model. In contrast, in similar settings $G-F$ is strictly decreasing at a unique finite equilibrium point. While not all settings will behave as nicely as our Gaussian example, under reasonable behavior in the tails we expect existence of a stable nontrivial equilibria, while the uninformative equilibria remain unstable. We discuss such generalizations in § B.
3.3 Gaussian Model
We revisit the Gaussian setting introduced in § 2.4 under the DG mechanism. First note that the same equilibria occur under DG as in OA.
Corollary 3.
[…]
Proof.
Existence of equilibria at $\tau=\pm\infty$ immediately follows from Proposition 3.
Now, note that $P(\tau;x)$ (Equation (7)) is strictly decreasing and continuous so that Theorem 3 applies. By Equation (8), $G(0)=F(0)=\Phi(0)$, so $\tau=0$ is an equilibrium under DG. Moreover, there could not be another finite equilibrium unless the coefficients of $x$ in Equation (8) were equal. Letting $c_{F}=\frac{\sqrt{\rho}}{a}$ and $c_{G}=\frac{(1-\rho)}{b\sqrt{1+\rho}}$ be these coefficients, and substituting $a^{2}=\frac{\rho}{1-\rho}b^{2}$, we have
$\left(\frac{c_{F}}{c_{G}}\right)^{2}=\frac{\rho b^{2}(1+\rho)}{a^{2}(1-\rho)^{2}}=\frac{\rho b^{2}(1+\rho)}{\frac{\rho}{1-\rho}b^{2}(1-\rho)^{2}}=\frac{1+\rho}{1-\rho}>1~{}.$ (13)
Uniqueness of $\tau=0$ as a finite threshold equilibrium follows. ∎
Now, consider the stability of these equilibria.
Corollary 4.
[…]
Proof.
As depicted in Figure 2, $G(x)-F(x)$ is strictly decreasing at 00. Formally, $G^{\prime}(x)-F^{\prime}(x)=c_{G}\phi(c_{G}x)-c_{F}\phi(c_{F}x)$, so $G^{\prime}(0)-F^{\prime}(0)=(c_{G}-c_{F})\phi(0)<0$ (since $c_{G}<c_{F}$ by (13)). Thus, by Theorem 4, the equilibrium at 00 is stable. Since $\tau=0$ is stable, the equilibria $\tau=\pm\infty$ are unstable by topological necessity. ∎
Since $\tau=0$ is the only stable equilibrium, it is in fact globally attracting: for any finite starting threshold $\tau(0)$, we have $\lim_{t\to\infty}\tau(t)=0$. In contrast to OA, this means that the behavior of DG in the Gaussian case is quite robust: unless the initial threshold starts exactly at an uninformative equilibrium it inevitably converges toward an informative one where $H$ and $L$ each get reported half the time. In this sense, DG behaves in a way one might hope based on its traditional analysis: it discourages uninformative equilibria.
While robust, this equilibrium is also inflexible: the mechanism designer cannot specify a desired threshold, and must settle for $\tau=0$. Thus the semantics of $H$ and $L$ relative to the underlying signal are predetermined, as half of the tasks must be classified each way. In peer grading, for example, half the assignments would be labeled unsatisfactory, which might not be the desired outcome. [^4] See § 7 for futher discussion. We expect this general phenomenon, that there is a unique equilibrium threshold for DG which is determined by the underlying signal distribution, extends beyond the Gaussian case; see § 6.
4 Determinant-based Mutual Information (DMI) Mechanism
The Determinant-based Mutual Information (DMI) mechanism was introduced by Kong [2024, 2020]. It is a minimal multi-task mechanism with several strong guarantees, such as being dominantly truthful under consistent strategies (see below). In the case of binary reports, the simplest version of the DMI mechanism collects 4 reports from a pair of agents, and computes a score based on a measure of their mutual information. To construct an unbiased estimator for this mutual information measure, the pairs of reports are split into two groups, and the determinants of the count matrices associated to each group are multiplied together.
Formally, the mechanism is given as follows. Define $\mathbf{1}_{H}=(1,0)$ and $\mathbf{1}_{L}=(0,1)$ to be indicator vectors in $\mathbb{R}^{2}$. Then $M_{\textrm{DMI}}:\{L,H\}^{4}\times\{L,H\}^{4}\to\mathbb{R}$ is the function
$M_{\textrm{DMI}}(r_{1},\ldots,r_{4},r_{1}^{\prime},\ldots,r_{4}^{\prime})=\det M_{12}\det M_{34}~{},$
where $M_{ij}=\mathbf{1}_{r_{i}}\mathbf{1}_{r_{i}^{\prime}}^{T}+\mathbf{1}_{r_{j}}\mathbf{1}_{r_{j}^{\prime}}^{T}$ is the count matrix for tasks $i$ and $j$.
4.1 Equilibrium Characterization
Unlike DG, where we could compute the payment for a single task in isolation (albeit requiring multiple tasks from a different agent), DMI is inherently a multi-task mechanism. As in the original analysis of DMI, here we restrict to consistent strategies, where the same $\sigma$ is applied to each $X_{i}$. (See below for behavior beyond this case.) Even under this restriction, threshold equilibria are much more complex in DMI given the interactions between four different signals, making it unclear how to reason directly from the definition of a threshold equilibrium. Instead, we restrict all agents to playing threshold strategies. Let
$U_{i}(\sigma,\sigma^{\prime})=\mathop{\mathbb{E}}\left[M_{\textrm{DMI}}(\sigma(X_{1}),\ldots,\sigma(X_{4}),\sigma^{\prime}(X_{1}^{\prime}),\ldots,\sigma^{\prime}(X_{4}^{\prime}))\right].$ (14)
be the (ex-ante) expected utility for playing threshold strategy $\sigma_{i}$ given the other agent using $\sigma_{j}$.
Definition 2.
A threshold strategy $\sigma:\mathbb{R}\to\mathcal{R}$ is a equilibrium restricted to threshold strategies under DMI if for all threshold strategies $\hat{\sigma}:\mathbb{R}\to\mathcal{R}$, $U_{i}(\sigma,\sigma)\geq U_{i}(\hat{\sigma},\sigma)$.
This is a weaker condition than being a threshold equilibrium becacuse it rules out deviations to non-threshold strategies, an issue we will return to. However, by restricting to this space we can provide an equilibrium characterization in the same spirit as we did for OA and DG.
As described above, the DMI mechanism is designed as an unbiased estimator of the (squared) determinant-based mutual information $\textrm{DMI}(Y;Y^{\prime})^{2}=(\det U_{Y;Y^{\prime}})^{2}$, where $U_{Y,Y^{\prime}}$ is the joint distribution matrix of the $\{H,L\}$-valued random variables $Y,Y^{\prime}$, given by
$U_{Y,Y^{\prime}}=\begin{bmatrix}\Pr[Y=H,Y^{\prime}=H]&\Pr[Y=H,Y^{\prime}=L]\\ \Pr[Y=L,Y^{\prime}=H]&\Pr[Y=L,Y^{\prime}=L]\end{bmatrix}~{}.$
This unbiasedness is implied by Kong [2024, Claim 4.4], which we paraphrase for this binary-report context: for any $Y,Y^{\prime}$, there exists a constant $\alpha>0$ such that $\mathop{\mathbb{E}}\det\sum_{i\in S}\mathbf{1}_{Y}\mathbf{1}_{Y^{\prime}}^{T}=\alpha\det U_{Y,Y^{\prime}}$.
Back to our real-valued setting, recall that $(X_{i},X_{i}^{\prime})\sim\mathcal{D}$ independently. Given strategies $\sigma,\sigma^{\prime}$ with thresholds $\tau,\tau^{\prime}$, for some $\alpha^{\prime}>0$ we can rewrite Equation (14) as $U_{i}(\sigma,\sigma^{\prime})=\alpha^{\prime}h(\tau,\tau^{\prime})^{2}$ for
$h(\tau,\tau^{\prime})=\Pr[X>\tau,X^{\prime}>\tau^{\prime}]\Pr[X\leq\tau,X^{\prime}\leq\tau^{\prime}]-\Pr[X>\tau,X^{\prime}\leq\tau^{\prime}]\Pr[X\leq\tau,X^{\prime}>\tau^{\prime}]~{},$ (15)
where $(X,X^{\prime})\sim\mathcal{D}$. We are now in a position to analyze equilibria and dynamics.
Equilibrium results.
Given the definition of $h$ from Equation (15), we can show that uninformative strategies $\tau^{*}=\pm\infty$ remain equilibria in the DMI mechanism. The proof is immediate, as in both cases one of the two probabilities in each term of Equation (15) is zero independent of $\tau$.
Proposition 5.
$\tau^{*}=\pm\infty$ are both always equilibria restricted to threshold strategies under DG.
We can also provide necessary and sufficient conditions for a finite threshold equilibrium for DMI. The theorem relies on several regularity conditions that rule out corner cases where the first order condition is trivially satisfied for spurious reasons.
Theorem 5.
Let finite threshold $\tau$ be given, $h(x,\tau)$ be differentiable in $x$ at $x=\tau$, $h(\tau,\tau)\neq 0$, and $f(\tau)>0$, where $f$ is the continuous density of $F$. If $\tau$ is an equilibrium restricted to threshold strategies under the DMI mechanism then $G(\tau)=F(\tau)$. Conversely, if $G(\tau)=F(\tau)$ and $h(x,\tau)$ is non-negative and strictly single-peaked then $\tau$ is an equilibrium restricted to threshold strategies under the DMI mechanism.
Proof.
For necessity, assume that $\tau^{*}$ is an equilibrium restricted to threshold strategies, so $f(\tau,\tau^{*})$ is maximized at $\tau=\tau^{*}$. The first order condition for optimality is
00 0 $=\frac{d}{d\tau}\alpha^{\prime}h(\tau,\tau^{*})^{2}=2h(\tau,\tau^{*})\alpha^{\prime}f(\tau)\left(\Pr[X^{\prime}>\tau^{*}]-\Pr[X^{\prime}>\tau^{*}\mid X=\tau]\right)~{}$ (16)
By assumption $h(\tau^{*},\tau^{*})\neq 0$ and $f(\tau^{*})>0$. Thus $\Pr[X^{\prime}>\tau^{*}]-\Pr[X^{\prime}>\tau^{*}\mid X=\tau^{*}]=0$, or equivalently $G(\tau^{*})-F(\tau^{*})=0$
For sufficiency, assume $G(\tau)=F(\tau)$. By the above, $\tau$ satisfies the first order condition for equilibrium. If $h(x,\tau)$ is non-negative and strictly single-peaked as a function of $x$ then $U_{i}$ is also strictly single-peaked as a function of $x$ (i.e. it is strictly increasing and then strictly decreasing), so the unique point satisfying the FOC is the global maximum. ∎
Theorem 5 shows that DMI has the same necessary condition for equilibrium as DG. So, modulo the slightly different sufficient condition for equilibrium, its equilibria can be found at the same thresholds: those where $G(\tau)=F(\tau)$. In particular our result about the dynamics (Theorem 4) immediately applies with the appropriate sufficient condition, and the Gaussian case has the same equilibra with the same stability.
Beyond consistent threshold strategies.
The above analysis is restricted to our weaker notion of equilibrium in the space of consistent threshold strategies. It is unclear whether these remain equilibria if agents can use any consistent strategy, or if agents can use any four threshold strategies. Even more broadly, one could consider joint-task strategies $\sigma:\mathbb{R}^{4}\to\{H,L\}^{4}$, which map all four signals simultaneously to all four reports. While not shown in Kong [2020, 2024], in the binary report setting the DMI mechanism turns out to satisfy the even stronger property that truthfulness is an equilibrium among all joint-task strategies. We prove this claim in Theorem 9 in § C. We then show that truthfulness is not an equilibrium for real-valued reports: in many cases the best response flips one or two of the truthful reports to increase the chance of a nonzero payoff in $M_{\textrm{DMI}}$.
5 Robust Bayesian Truth Serum
We now study the Robust Bayesian Truth Serum (RBTS) mechanism proposed by Witkowski and Parkes [2012]. RBTS is inspired by the Bayesian Truth Serum (BTS) [Prelec, 2004]. However, while incentive alignment under BTS requires an arbitrarily large number of agents, RBTS guarantees truthfulness for $n\geq 3$ agents.
RBTS is a non-minimal mechanism: each agent $i$ submits both an information report $r_{i}\in\{L,H\}$ and also a prediction report $p_{i}\in[0,1]$. The prediction report represents agent $i$’s belief about the frequency of $H$ reports from all agents. The mechanism incentivizes prediction reports using the Brier score $S(p,r)=2p\mathbf{1}[r=H]-p^{2}$ (also known as the quadratic score), which is an example of a strictly proper scoring rule [Brier, 1950, Gneiting and Raftery, 2007]. RBTS randomly picks a reference agent $j$ and peer agent $k$, and pays agent $i$
$M_{\textrm{RBTS}}((r_{i},p_{i}),(r_{j},p_{j}),(r_{k},p_{k}))=S(p_{j}+\delta t(r_{i}),r_{k})+S(p_{i},r_{k}),$ (17)
where $t(L)=-1$ and $t(H)=1$, and $\delta>0$ is chosen by the mechanism.
We refer to the first term in Equation 17 as the information score, and the second as the prediction score. Under binary signals and reports, truthfulness over both the information and prediction report forms a strict Bayes-Nash equilibrium.
5.1 Equilibrium Characterization
If we consider a threshold strategy $\tau$ in our real-valued signal model, an agent with signal $x_{i}$ has belief $p_{i}(x_{i})=\Pr[X^{\prime}>\tau\mid X=x_{i}]=1-P(\tau;x_{i})$. Since the Brier score $S$ is proper, it immediately follows that agents will report $p_{i}$ truthfully in a symmetric threshold Bayes-Nash equilibrium. Thus it is WLOG that we can study incentives in threshold equilibria under only the information score, $M_{\textrm{RBTS}}(r_{i},p_{j},r_{k})=S(p_{j}+\delta t(r_{i}),r_{k})$. Moreover, we can operate under the assumption that the prediction report $p_{j}$ is truthful, i.e. $p_{j}=p_{j}(x_{j})=1-P(\tau;x_{j})$ for agent $j$’s true signal $x_{j}$.
Assume an agent’s signal $x$ satisfies $x\leq\tau$. Then a threshold equilibrium $\tau$ must satisfy
$\mathop{\mathbb{E}}_{x_{j},x_{k}\sim\beta(x)}\left[S\left(p_{j}(x_{j})-\delta,\sigma(x_{k})\right)\right]$ $\geq\mathop{\mathbb{E}}_{x_{j},x_{k}\sim\beta(x)}\left[S\left(p_{j}(x_{j})+\delta,\sigma(x_{k})\right)\right]$
$\mathop{\mathbb{E}}\left[2(p_{j}(x_{j})-\delta)\mathbf{1}[\sigma(x_{k})=H]-(p_{j}(x_{j})-\delta)^{2}\right]$ $\geq\mathop{\mathbb{E}}\left[2(p_{j}(x_{j})+\delta)\mathbf{1}[\sigma(x_{k})=H]-(p_{j}(x_{j})+\delta)^{2}\right]$
$4\delta\mathop{\mathbb{E}}\left[p_{j}(x_{j})-\mathbf{1}[\sigma(x_{k})=H]\right]$ $\geq 0$
$\Pr[X^{\prime}\leq\tau\mid X=x]$ $\geq\mathop{\mathbb{E}}_{x_{j}\sim\beta(x)}\left[\Pr[X^{\prime}\leq\tau\mid X_{j}=x_{j}]\right]$
where the last line follows from plugging in $p_{j}(x_{j})=1-\Pr[X^{\prime}\leq\tau\mid X_{j}=x_{j}]$, and noting that $\mathop{\mathbb{E}}_{x_{j},x_{k}\sim\beta(x)}\mathbf{1}[\sigma(x_{k})=H]=1-\Pr[X^{\prime}\leq\tau\mid X=x]$. One can show a symmetric condition if $i$’s signal satisfies $x>\tau$, so that all symmetric Bayes-Nash threshold equilibria $\tau$ are characterized by the following conditions:
$\forall x\leq\tau,P(\tau;x)$ $\geq\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}P(\tau;x^{\prime}),$ (18)
$\forall x>\tau,P(\tau;x)$ $\leq\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}P(\tau;x^{\prime}).$ (19)
Therefore, unlike the DG mechanism, in RBTS an agent’s actions in equilibrium depend on the expected conditional probability distribution. We can immediately see that the uninformative thresholds $\tau=\pm\infty$ still always appear as equilibria:
Proposition 6.
$\tau^{*}=\pm\infty$ are threshold equilibria under RBTS.
Proof.
Let $\tau^{*}=\infty$. Then for all signals $x,x^{\prime}\in\mathbb{R}$, $P(\tau^{*};x)=1$ and $P(\tau^{*};x^{\prime})=1$, so that each side of Equation 18 always evaluates to 1 and the condition for $x<\tau^{*}$ is satisfied. The condition of Equation 19 is vacuous. An analagous argument holds for $\tau^{*}=-\infty$. ∎
We also observe necessary and sufficient conditions for non-infinite equilibria $\tau$, relative to the function $G$. Let $Q(x)=\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}P(x;x^{\prime})$, and let $\bar{P}(\tau;x)=\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}P(\tau;x^{\prime})$ be the expected conditional probability of another agent reporting $L$ over signal $x$.
Theorem 6.
Let a finite threshold $\tau$ be given and $P(\tau;x)$ be continuous over $x$; also assume $\beta(x)$ is continuous over $x$. If $\tau\in\mathbb{R}$ is threshold equilibrium under the RBTS mechanism then $Q(\tau)=G(\tau)$. Conversely, if $Q(\tau)=G(\tau)$ and $P(\tau;x)-\bar{P}(\tau;x)$ has a single crossing of 00 from positive to negative then $\tau$ is a threshold equilibrium under the RBTS mechanism.
Proof.
For necessity, first note that since $\beta(x)$ and $P(\tau;x)$ are continuous over $x$, $\bar{P}(\tau;x)=\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}P(\tau;x^{\prime})$ is also continuous over $x$ by an application of the Dominated Convergence Theorem. Let $f(\tau;x)=P(\tau;x)-\bar{P}(\tau;x)$. Since $f(\tau;x)$ is continuous, we must have $\lim_{x\to\tau^{+}}f(\tau;x)=\lim_{x\to\tau^{-}}f(\tau;x)=f(\tau;\tau).$ Further, $\lim_{x\to\tau^{-}}f(\tau;x)\geq 0$ and $\lim_{x\to\tau^{+}}f(\tau;x)\leq 0$ by Conditions 18 and 19, so $f(\tau;\tau)=0=G(\tau)-Q(\tau)$.
For sufficiency, take some $x>\tau$. By single crossing, $P(\tau;x)\leq\bar{P}(\tau;x)$. Similarly, if $x\leq\tau$, $P(\tau;x)\geq\bar{P}(\tau;x)$. This establishes Equations (18) and (19), so $\tau$ is a threshold equilibrium. ∎
While the exact crossing condition has changed from $G(\tau)=F(\tau)$ to $G(\tau)=Q(\tau)$, Theorem 6 paints a similar picture to what we saw for DG. Best responses arew now more complex to compute, so we do not treat the general dynamics formally, but we can provide a comparison of their behavior in the Gaussian case.
5.2 Gaussian Model
We return to our Normal example, where each agent receives a noisy version of a Gaussian signal. Leaving the technical details to § D, we first note that the same equilibria occur as the previous mechanisms.
Corollary 5.
[…]
We also derive the exact form of the best response $\hat{\tau}=c(\rho)\tau$ in § D. Now, note that for any $\tau$, $\hat{\tau}=c(\rho)\tau$ for $c(\rho)\in(0,1)$. Under our dynamics, $\dot{\tau}=\hat{\tau}-\tau.$ If $\tau>0$, then $\hat{\tau}<\tau$, while if $\tau<0$, $\hat{\tau}>\tau$. It follows that $\tau=0$ is stable, while by topological necessity the uninformative equilibria are unstable.
While the equilibria and dynamics behave the same in RBTS as in DG for the Gaussian model, we can ask if there are differences in the convergence rate to stable point $\tau=0$. In fact, we can observe that the best response in both settings is a linear function of $\tau$ with coefficients $m_{\textrm{DG}}$, $m_{\textrm{RBTS}}$ respectively (see § D for a formal derivation). It turns out that in any setting of parameters $a$ and $b$, $m_{\textrm{RBTS}}>m_{\textrm{DG}}$, so that if both mechanisms begin with the same starting threshold $\tau(0)$, the stable equilibrium at 00 is reached faster in DG than RBTS. In a peer grading setting, then, graders will more quickly converge to a half-good half-bad quality consensus than RBTS graders, so that the designer may have more time to react and reset expectations in the latter case.
6 Experiments
We now aim to numerically explore existence and dynamics of equilibria across mechanisms in settings other than the noisy Gaussian model. We consider two natural departures from the “nice” properties of Gaussians, skewness and multimodality.
6.1 Skewness
Skewed models of information capture settings where there is some asymmetry in the tasks. To study such settings, we adapt the same noisy signal model as our running Gaussian example, but with a base distribution which is skewed. Formally, each agent $i$ receives a signal $X_{i}$ such that $X_{i}=Z+Z_{i}$, where $Z$ is the skewed Normal distribution with mean 0 and variance 1, and $Z_{i}\sim N(0,1)$. $Z$ has another parameter, $\alpha$, which controls the skewness of the distribution. The magnitude of $\alpha$ increases the magnitude of skewness; moreover, $\alpha<0$ leads to a left-skewed distribution, and $\alpha>0$ a right-skewed distribution. We study how threshold equilibria move as a function of the parameter $\alpha$.
As depicted in Figure 4, we find a unique equilibrium under both OA and DG which shifts toward the longer tail, i.e. the thresholds move from negative to positive as the skewness moves from left to right (and we recover the $\tau=0$ equilibrium in the original Gaussian model when $\alpha=0$). Thus we expect when there is some asymmetry in the distribution of signals, agents in DG will settle at a threshold that balances out the asymmetry in signal mass. While we observe the same high-level dynamics for OA, these equilibria are unstable and so we still expect agents will move toward an uninformative consensus. For DMI, meanwhile, we note that the same equilibria will occur as DG by our theoretical results. Moreover, given the behavior observed under the Gaussian, we expect the stable equilibria in RBTS to share the same pattern as DG in tracking the skewness of the distribution, but at a rate in between OA and DG.
6.2 Multimodality
We now explore the effects that a distribution with multiple peaks has on the equilibria that occur under our signal model. Consider a classic mixed Gaussian setting where each agent’s signal $X_{i}$ is received i.i.d. from one of $K$ Gaussian components, each with mean $\mu_{k}$ and variance $\sigma_{k}^{2}$. There is an underlying parameter $z\sim\text{Categorical}(\pi)$ for $\pi\in\Delta^{K}$ that determines which Gaussian component all the signals come from. For the purposes of our experiments, we fix $\pi$ to be the uniform distribution over the $K$ components. We numerically calculate the functions $F$, $G$, and $Q$, and check for the sufficient and necessary conditions of our theorems, to find equilibria and their stability in this setting. We set the precision of each component to be the same, and then vary this precision parameter to generate bifurcation graphs and study how these equilibria evolve for $K=2$ and $3$ (see § E for $K=4$).
Figure 3: Equilibria in OA and DG when signals are noisy versions of a skewed Normal, across different values of the skewness parameter $\alpha$.
Figure 4: $F$, $G$, and $Q$ over signals $x$ in the two-component Gaussian mixture setting with means $-2,2$ and precision 2.
Figure 5: Bifurcation diagrams for two- and three-component Gaussian mixtures. (Top) Two-component case: OA (left) exhibits three equilibria at high precision, while DG (right) maintains a single stable equilibrium at $\tau=0$. (Bottom) Three-component case: OA (left) has five equilibria, with two stable at $\tau=\pm 1$; DG (right) has three equilibria, with stability reversed from the two-component OA case.
We first observe that based on our experiments, the number of equilibria increases by two each time we add a Gaussian component (Figure 5). But OA begins with three equilibria for large enough precision under the two-component model, while DG begins with one. Since a unique equilibrium in DG is stable, this is a testament to DG being robust to uninformative equilibria. However, for a high enough separation between the two modes, stable equilibria emerge (at $\tau=0$ in the three component case) in OA, and the area of starting conditions that lead to this equilibrium increases. One can view this phenomenon as a reflection of the original binary signal model: with enough separation between the two modes of the richer signal space, we essentially return to a setting where Output Agreement is useful because the “signal” is effectively which mode has been chosen.
We can view equilibria in the snapshot of a specific precision value in Figure 4. Here, the intersection of $G$ with 1/2 corresponds to equilibria in OA, with $F$ corresponds to equilibria in DG (and DMI), and with $Q$ corresponds to equilibria in RBTS. The function $Q$ closely tracks $G$ for high enough precision values. We therefore find that the same stability and number of equilibria occur in RBTS as in DG. However, we note that bifurcation (the increase in equilibria) occurs at lower precision values for RBTS (see § E for visualization). From a design perspective, a lower bifurcation point in RBTS is nice because there is a wider set of signal distributions allowing for some choice over stable thresholds. Meanwhile, our theoretical results show that equilibria restricted to the threshold space in DMI correspond to those in DG, so that we would expect the trends we observe over precision to transfer directly from DG to DMI.
In general, then, our simulations show that multiple threshold equilibria occur across mechanisms with numbers increasing as modes separate, but similar patterns of stability emerge. Moreover, it is even more clear how different behavior emerges under a richer signal model. One could imagine two different real-valued distributions, the first unimodal as in a peer grading setting and the second bimodal as in a data labeling setting, with the same collapsed binary signal model. While previous results on the collapsed space concludes the same behavior would emerge in both settings, our results show that the number and stability of equilibria will be significantly different.
7 Discussion
We study behavior in several peer prediction mechanisms when signals are real-valued and reports are binary. While all these mechanisms have various truthfulness guarantees in the binary signal setting, in our model we find that the notion of truthfulness breaks. Specifically, when agents play according to thresholds which partition the signal space into “high” or “low” regions, under dynamics we find equilibria rarely stay at the initial threshold a mechanism designer sets. Instead, for several mechanisms (Dasgupta–Ghosh, Determinant Mutual Information, and the Robust Bayesian Truth Serum) agents will often stabilize at a different threshold according to the specific payment scheme. Moreover, in Output Agreement we find that uninformative equilibria are often stable, so that in many information structure settings agents will always shift toward an uninformative consensus. We show how such equilibria play out in a natural, noisy Gaussian signal setting and extend our analysis to experiments with multimodal and skewed distributions.
Our results imply that the standard modeling assumptions in peer prediction literature miss important nuances, and these holes can propagate to real changes in how we expect agents to behave in practice. In particular, models of peer prediction have often taken as given an arbitrary joint distribution $P$ over pairs of signals drawn from a finite set. This suggests that we should be able to choose an arbitrary meaning for each signal and then peer prediction should “just work” for the resulting distribution. In contrast, our results show that peer prediction is substantially more inflexible in practice. Whenever the meaning of signals has endogeneity to it there will typically be only a small number of plausible equilibria in this richer space.
Adding flexibility.
We saw in the Gaussian case that a 50/50 report distribution was a natural outcome, which might be unsuitable for some applications. For example in peer grading or norm enforcement [Alechina et al., 2017], one might want the final decision for most tasks to be “accept”. So what should a mechanism designer do? Since peer prediction naturally relies on multiple reports about each task, one option is to use a decision rule, which combines multiple reports on a task in a way that is aware of this limitation. In peer grading an assignment might be satisfactory if any peer grader reports $H$. Doing so would allow more flexibility on the final decisions, e.g. increasing the fraction of satisfactory assignments, even under fixed semantics of $H$ and $L$.
An alternative is to seek to design new families of peer prediction mechanisms that allow control over where the equilibrium thresholds occur. The fact that our characterizations yielded different conditions among our four mechanisms suggests at least some scope for this, although the the fact that in the Gaussiance case these conditions coincided also points to potential limitations. Nevertheless, this seems an interesting direction for future work.
Beyond binary reports.
[…]
Effort.
Studying effort would be natural in our model. For example, one could model effort in the Gaussian model as giving the agent $aZ+bZ_{i}$ where $b$ is decreasing in their exerted effort; in the limit, they obtain the “ground truth” and need only contend with the noisy reports of others. One interesting interplay with our dynamics is that low-effort agents seem most likely to move the threshold, as high-effort agents will have more reason to respect the current one; e.g. when all agents are high-effort, they all coordinate perfectly on the correct report.
Further broadening the model.
We have peeled back the peer prediction abstraction by one layer, from binary signals to real-valued ones. This already substantially changed the behavior of mechanisms. Yet even this is quite a stylized abstraction of the information available. Another missing piece is the fact that in many settings some deterministic context, like the content of the essay being graded, is shared by all looking at the same task. From this “shared signal” comes the well-known problem of spurious correlations where in principle participants may choose strategies that depend on inessential details of the context, like the first word of the essay, to correlate their reports without correlating with the ground truth, all while getting optimal payoffs in multi-task mechansisms. While there is work seeking to deal with heterogeneity among agent behaviors [Agarwal et al., 2020], such behavior seems impossible to fully rule out. But even if such pathological equilibra seem unlikely in systems with many well-meaning participants, there is often still ample metadata about tasks such as categories or difficulties that naturally influence participants’ process of translating their actual experience into a “signal.” Designing mechanisms that are robust in the face of such richness is a further challenge.
Acknowledgments
This paper is based on initial results when the first two authors were working with the Oinc cryptocurrency team. We thank Tim Roughgarden, and members of Oinc and Kleros, for useful discussions.
Appendix A Output Agreement
A.1 Equilibrium Characterization Generalization
In general, we can identify where finite equilibria exist in OA based on the extreme behavior of the function $G$.
Condition 1.
Let $I=[a,b]\subset\mathbb{R}$ be an interval. Upon seeing signal $x\leq a$, then an agent believes with probability less than 1/2 that another signal will be less than $x$. Meanwhile, upon seeing signal $x\geq b$, then an agent believes with probability greater than 1/2 that another signal will be less than $x$. Formally, $x\leq a\implies P(x^{\prime}\leq x\mid x)<1/2$ and $x\geq b\implies P(x^{\prime}\leq x\mid x)>1/2$.
Theorem 7.
Let the agent signal structure satisfy Condition 1, with $\Pr[X^{\prime}\leq\tau\mid X=x]$ monotone decreasing and continuous in $x$, and $G$ continuous. Then there exists an equilibrium $\tau^{*}\in I$.
Proof.
Condition 1 implies $G(a)<1/2$ and $G(b)>1/2$. Since $G$ is continuous, we can apply the Intermediate Value Theorem to conclude there exists a point $\tau^{*}\in[a,b]$ such that $G(\tau^{*})=1/2$. Then $\tau^{*}$ is an equilibrium by Theorem 1. ∎
Note in particular that if $G(x)$ is monotone increasing and crosses 1/2, Condition 1 is satisfied for the single point interval $\tau$ where $G(\tau)=1/2$ and we have a unique equilibrium.
Proposition 7.
Let the agent signal structure satisfy Condition 1, with $\Pr[X^{\prime}\leq\tau\mid X=x]$ monotone decreasing and continuous in $x$, and $G$ continuous. Let $m$ be the median of $F$, and consider running the OA mechanisim. Then if $\Pr[X^{\prime}\leq m\mid m]>1/2$, there exists an equilibrium $\tau^{*}$ such that $\tau^{*}<m$. Similarly, if $\Pr[X^{\prime}\leq m\mid m]<1/2$, there exists an equilibrium $\tau^{*}$ such that $\tau^{*}>m$.
Proof.
First, by Theorem 7, we know at least one equilibrium $\tau^{*}$ exists in the interval $I$. Now assume that $\Pr[X^{\prime}\leq m\mid m]=G(m)>1/2$. It follows under Condition 1 that either $m>b$ or $m\in I$. If $m>b$, then by Theorem 7 there exists an equilibrium $\tau^{*}$ in the interval $I$; $\tau^{*}\leq b<m$, so the result follows. If $m\in I$, then $m\geq a$. Since $G$ is continuous, $G(a)<1/2$, and $G(m)>1/2$, by the Intermediate Value Theorem there exists a point $\tau^{*}\in[a,m]$ such that $G(\tau^{*})=1/2$. By our characterization in Theorem 3, $\tau^{*}$ is an equilibrium.
A symmetric argument holds for when $\Pr[X^{\prime}\leq m\mid m]<1/2$. ∎
A.2 Dynamics Generalization
We can observe instability of equilibria under Condition 1.
Proposition 8.
[…]
Proof.
We know already from Theorem 7 that at least one equilibrium exists in the interval $I$. Note first that since $G(a)<1/2$ and $G(b)>1/2$, by continuity and differentiability $G$ must cross 1/2 at some point $\tau_{0}$ with $G^{\prime}(\tau_{0})>0$. Assume not: then $G(x)\leq 1/2$ for all $x$, which contradicts the fact that $G(x)>1/2$ for $x\geq b$. It follows by Theorem 2 that $\tau_{0}$ is an unstable equilibrium.
  1. 1.
    Now, WLOG we pick the equilibrium $\tau_{1}$ which is closest to $a$. By Condition 1 and since $G(x)$ is continuous, $G(x)<1/2$ for $x<\tau_{1}$. Then it must follow that $G^{\prime}(\tau_{1})\geq 0$. Thus $\tau_{1}$ is unstable, at least on the left (if $G^{\prime}(\tau_{1})=0$, then $\tau_{1}$ is unstable on the left and stable on the right).
  2. 2.
    Pick the equilibrium $\tau_{2}$ which is closest to $b$ (this could be the same as the $\tau_{1}$ we chose in the previous step). By Condition 2 and since $G(x)$ is continuous, $G(x)>1/2$ for $x>\tau_{2}$. Then it must follow that $G^{\prime}(\tau_{2})\geq 0$. Thus $\tau_{2}$ is unstable, at least on the right (if $G^{\prime}(\tau_{2})=0$, then $\tau_{2}$ is unstable on the right and stable on the left).
[…]
In most reasonable settings, one would expect $G(\tau)$ to behave according to Condition 1. Specifically, as a signal $x$ becomes increasingly small, the probability another agent receives a signal smaller than $x$ should approach 0. Meanwhile, as a signal $x$ grows large, the probability another agent receives a signal smaller than $x$ should approach 1. Proposition 8 thus suggests that when information corresponds predictably and monotonically to quality of an essay or task, over time agents will naturally move toward uninformative equilibria. Intuitively, these dynamics make sense: an agent who thinks others are submitting more “high” reports will shift their threshold down to match, reinforcing larger and larger thresholds over time.
Appendix B Dasgupta-Ghosh
B.1 Omitted Proofs
Proof of Proposition 3.
Let $\tau^{*}=\infty$. Then for all signals $x$ such that $x\leq\tau^{*}$, $P(\tau^{*};x)=1\geq\Pr[X^{\prime}\leq\infty]$ (since any $X^{\prime}\in\mathbb{R}$ satisfies $X^{\prime}<\tau^{*}$ with probability one.) Meanwhile, Statement 12 is vacuous since no signal $x\in\mathbb{R}$ satisfies $x>\infty$. A similar argument follows for $\tau^{*}=-\infty$: for all signals $x>\tau^{*}$, $P(\tau^{*};x)=0\leq\Pr[X^{\prime}\leq-\infty]$, while Statement 11 is vacuous since no signal $x\in\mathbb{R}$ satisfies $x<-\infty$. ∎
Proof of Theorem 3.
For necessity, assume that $\tau$ is a threshold equilibrium, so Equations (11) and (12) hold. Since $P(\tau;x)$ is continuous, we must have $\lim_{x\to\tau^{+}}P(\tau;x)=\lim_{x\to\tau^{-}}P(\tau;x)=P(\tau;\tau).$ Further, $\lim_{x\to\tau^{+}}P(\tau;x)\leq\pi_{L}$ and $\lim_{x\to\tau^{-}}P(\tau;x)\geq\pi_{L}$, so $G(\tau)=P(\tau;\tau)=\pi_{L}=F(\tau)$.
For sufficiency, assume $G(\tau)=F(\tau)$. Equivalently $P(\tau;\tau)-\pi_{L}=0$, so (a) implies (b) and the single crossing is at $\tau$. Thus assume (b) holds and take some $x>\tau$. By single crossing, $P(\tau;x)\leq P(\tau;\tau)=G(\tau)=F(\tau)=\pi_{L}$. Similarly, if $x\leq\tau$, $P(\tau;x)\geq P(\tau;\tau)=\pi_{L}$. This establishes Equations (11) and (12), so $\tau$ is a threshold equilibrum. ∎
Proof of Proposition 4.
If $\tau$ is the current threshold strategy that all other agents are following, an agent who receives signal $x$ will best respond with $H$ if $P(\tau;x)\leq F(\tau)$, and $L$ otherwise. Since $P(\tau;x)$ is strictly decreasing and continuous over $x$, there will be a unique point $\hat{\tau}$ such that $P(\tau;\hat{\tau})=F(\tau)$, with $P(\tau;x)\geq F(\tau)$ for $x\leq\hat{\tau}$ and $P(\tau;x)\leq F(\tau)$ for $x>\hat{\tau}$. This threshold $\hat{\tau}$ thus corresponds to the best response. ∎
Proof of Theorem 4.
Consider the dynamics at time step $t$, with the current threshold $\tau(t)$. Since we are considering a fixed time $t$, we refer to $\tau(t)$ as $\tau$ throughout the proof.
We consider what happens when an agent receives a signal exactly at $\tau$. There are three cases to consider. In the first, $G(\tau)=F(\tau)$; then the expected utility of reporting $L$ is $\Pr[X^{\prime}\leq\tau\mid X=\tau]-\Pr[X^{\prime}\leq\tau]=0$, while the expected utility of reporting $H$ is $\Pr[X^{\prime}>\tau\mid X=\tau]-\Pr[X^{\prime}>\tau]=0$; thus the agent is indifferent between reporting $H$ or $L$. Moreover, the best response $\hat{\tau}$ is exactly $\tau$. It follows the system is at an equilibrium since $\dot{\tau}=\hat{\tau}-\tau=0$.
In the second case, $G(\tau)>F(\tau)$. Then reporting $L$ is strictly preferred to reporting $H$. Since $P(\tau;\hat{\tau})=F(\tau)<G(\tau)=P(\tau;\tau)$, by decreasing monotonicity we have $\hat{\tau}>\tau$. Thus $\dot{\tau}=\hat{\tau}-\tau>0$. In the third case, $G(\tau)<F(\tau)$. Then reporting $H$ is strictly preferred to reporting $L$. Since $P(\tau;\hat{\tau})=F(\tau)>G(\tau)=P(\tau;\tau)$, by decreasing monotonicity we have $\tau>\hat{\tau}$. Thus $\dot{\tau}=\hat{\tau}-\tau<0$.
If $G(\tau)-F(\tau)$ is strictly decreasing at $\tau^{*}$, then at a sufficiently close perturbed point $\tau$ to the left of $\tau^{*}$ we have $G(\tau)>F(\tau)$ and $\dot{\tau}>0$; and at a perturbed point $\tau$ to the right of $\tau^{*}$ we have $G(\tau)<F(\tau)$ and $\dot{\tau}<0$. Stability follows. The same logic implies instability in the strictly increasing case. ∎
B.2 Equilibrium Characterization Generalization
As in OA, while we do not expect all settings to behave as symmetrically as the Gaussian model in Section 2.4, we can still characterize existence and location of finite threshold equilibria based on the shapes of the functions $F$ and $G$.
Condition 2.
Let $I=[a,b]\subset\mathbb{R}$ be an interval. Upon seeing signal $x\leq a$, then relative to the prior, an agent believes it more likely that another signal will be less than $x$. Meanwhile, upon seeing signal $x\geq b$, then relative to the prior, an agent believes it more likely that another signal will be greater than $x$. Formally, $x\leq a\implies P(x^{\prime}\leq x\mid x)>P(x^{\prime}\leq x)$ and $x\geq b\implies P(x^{\prime}\geq x\mid x)>P(x^{\prime}\geq x)$.
Theorem 8.
Let the agent signal structure satisfy Condition 2, with $\Pr[X^{\prime}\leq\tau\mid X=x]$ monotone decreasing and continuous in $x$, and $F$ and $G$ continuous. Then there exists an equilibrium $\tau^{*}\in I$.
Proof.
Take the function $H(x)=G(x)-F(x)$. Then Condition 2 implies $H(a)=G(a)-F(a)>0$ and $H(b)=G(b)-F(b)<0$. Since $G$ and $F$ are continuous, $H$ is continuous; thus we can apply the Intermediate Value Theorem to conclude there exists a point $\tau^{*}\in[a,b]$ such that $H(\tau^{*})=0$. Then $\tau^{*}$ is an equilibrium by Theorem 3. ∎
If Condition 2 holds and $F$ is symmetric about a point (its median) in the interval $I$, as in the Gaussian model in Section 2.4, it is reasonable to expect that $G$ is symmetric about the same point and an equilibrium occurs at the median. However, we can also characterize how equilibria change relative to the median in skewed settings.
Proposition 9.
Let the agent signal structure satisfy Condition 2, with $\Pr[X^{\prime}\leq\tau\mid X=x]$ monotone decreasing and continuous in $x$, and $G$ and $F$ continuous. Let $m$ be the median of $F$, and consider running the DG mechanism. Then if $\Pr[X^{\prime}\leq m\mid m]>1/2$, there exists an equilibrium $\tau^{*}$ such that $\tau^{*}>m$. Similarly, if $\Pr[X^{\prime}\leq m\mid m]<1/2$, there exists an equilibrium $\tau^{*}$ such that $\tau^{*}<m$.
Proof.
First, by Theorem 8, we know at least one equilibrium $\tau^{*}$ exists in the interval $I$. Now assume that $\Pr[X^{\prime}\leq m\mid m]>1/2$. In other words, since $F(m)=1/2$, $G(m)>F(m)$. It follows under Condition 2 that either $m<a$ or $m\in I$. If $m<a$, then by Theorem 8 there exists an equilibrium $\tau^{*}$ in the interval $I$; $\tau^{*}\geq a>m$, so the result follows. Otherwise, consider the case where $m\in I$. Then define $H(x)=G(x)-F(x)$. $H(m)>0$, while under Condition 2, $H(b)<0$. By the Intermediate Value Theorem, since $H$ is continuous, there exists a point $\tau^{*}\in[m,b]$ such that $H(\tau^{*})=0$. By our characterization in Theorem 3, $\tau^{*}$ is an equilibrium.
A symmetric argument holds for when $\Pr[X^{\prime}\leq m\mid m]<1/2$. ∎
In unimodal, continuous settings, we can often diagnose skewness of a distribution based on the difference between its mean and median. In a right-skewed distribution, it is often the case that the bulk of high probability outcomes lie below the median, so that conditioning on the median essentially amplifies the impact of these outcomes. We would therefore expect $\Pr[X^{\prime}\leq m\mid m]$ to increase relative to 1/2. By Proposition 9, equilibria drift to the right of $F$’s median. A similar argument holds for left-skewed distributions: there is more high probability mass concentrated to the right of the median, so conditioning on the median should decrease the probability of a signal being below the median. Thus we would expect by Proposition 9 that when $F$ is left-skewed, equilibria drift to the left of $F$’s median.
B.3 Dynamics Generalization
Under reasonable behavior in the tails, we expect existence of a stable nontrivial equilibria, while the uninformative equilibria remain unstable.
Proposition 10.
[…]
Proof.
We know already from Theorem 8 that at least one equilibrium exists in the interval $I$. Again, we let $H(x)=G(x)-F(x)$. Note first that since $H(a)>0$ and $H(b)<0$, by continuity and differentiability $H$ must cross 0 at some point $\tau_{0}$ with $H^{\prime}(\tau_{0})<0$. Assume not: then $H(x)\geq 0$ for all $x$, which contradicts the fact that $H(x)<0$ for $x\geq b$. It follows by Theorem 4 that $\tau_{0}$ is a stable equilibrium.
  1. 1.
    Now, WLOG we pick the equilibrium $\tau_{1}$ which is closest to $a$. By Condition 2 and since $H(x)$ is continuous, $H(x)>0$ for $x<\tau_{1}$. It follows that $H^{\prime}(\tau_{1})\leq 0$. Then $\tau_{1}$ is stable, at least on the left (if $H^{\prime}(\tau_{1})=0$, then $\tau_{1}$ is stable on the left and unstable on the right).
  2. 2.
    Pick the equilibrium $\tau_{2}$ which is closest to $b$ (this could be the same as the $\tau_{1}$ we chose in the previous step). By Condition 2 and since $H(x)$ is continuous, $H(x)<0$ for $x>\tau_{2}$. It follows taht $H^{\prime}(\tau_{2})\leq 0$. Then $\tau_{2}$ is stable, at least on the right (if $H^{\prime}(\tau_{2})=0$, then $\tau_{2}$ is stable on the right and unstable on the left).
[…]
In natural settings, we would expect Condition 2 to hold. Specifically, conditioning on a small enough signal, one would expect the probability of other small signals to increase relative to the prior. Conditioning on a large signal, one would expect decreased probability of much smaller signals and thus a smaller value of $G$ relative to the prior. Ultimately, then, when conditioning on a signal leads to local sensitivity, we would expect the dynamics in Proposition 10 to hold. In particular, note that Proposition 10 applies to the model in § 6, when the distributions are skewed. Even under complicated, multi-modal information structures, we expect tail behavior to often still satisfy Condition 2, so that there exists a stable equilibrium in the interval.
Appendix C Omitted Proofs for DMI
As discussed in Section 4, the behavior of DMI under joint-task strategies, which allow agents to map all 4 signals to all 4 reports simultaneously, differs between the traditional binary signal model and our real-valued signal model. In particular, Theorem 9 shows that truthfulness is a Bayes Nash among all joint-task strategies for the binary signal model, but below we show that this property fails for the real-valued signal model.
To get some intuition for these results, first consider the binary signal model and an agent who receives signals $H,L,L,L$. From the form of the mechanism, if they report truthfully, their score will be zero deterministically, since $M_{34}$ will have rank 1. Thus, they might consider deviating, say flipping the last $L$ to $H$. It turns out doing so keeps their expected score at zero, since the other agent is equally likely to align or misalign with their reports. The consistency of the other agent’s strategies is crucial for this last fact; given any slight difference in their strategy for tasks 3 and 4, the first agent would deviate from truthfulness.
It is precisely this sort of imbalance that easily arises in the real-valued signal model. Consider the Gaussian model with $a=b=1$ and $\rho=1/2$, at the equilibrium threshold $\tau=0$. Suppose the agent receives $x_{1}=10$, $x_{2}=-10$, $x_{3}=-10$, and $x_{4}=-1$. Reporting truthfully, $(H,L,L,L)$, would yield payoff 0 deterministically. But the agent may note that flipping the last $L$ to $H$ is “safe” in the sense that it is extremely unlikely that the other agent will differ in their reports on the first 3 tasks. So either the other agent reports $r_{4}=L$, in which case the payment is zero for both agents, or $r_{4}=H$, in which case both agents receive 1. As there is a reasonable chance of the latter case, about $0.2819$, that is roughly the expected score.
Thus, given real-valued signals $(10,-10,-10,-1)$, the optimal report is $(H,L,L,H)$, which is not “truthful” for the prescribed threshold $\tau=0$. Hence DMI is no longer truthful in this joint-task sense, even at the equilibrium threshold $\tau=0$.
C.1 Proof of Theorem 9
First, we prove that in the binary signal model, truthfulness is an equilibrium among all joint-task strotegies $\sigma:\{H,L\}^{T}\to\Delta(\{H,L\}^{T})$. Truthfulness remains a best response even when the other agent deviates from truthful, but crucially, still plays a consistent strategy, which here means strategies that (a) map each signal to a distribution over reports, and (b) use the same such map for every task.
Definition 3.
A strategy $\sigma:\{H,L\}^{T}\to\Delta(\{H,L\}^{T})$ is consistent if $\sigma(s_{1..T})(r_{1..T})=\prod_{t=1}^{T}\hat{\sigma}(s_{t})(r_{t})$ for some $\hat{\sigma}:\{H,L\}\to\Delta(\{H,L\})$.
For example, the truthful strategy $\sigma_{\text{true}}:s\mapsto\delta_{s}$, where $\delta_{s}$ is the point distribution on $s$, is a consistent strategy, with $\hat{\sigma}:s\mapsto\delta_{s}$ for $s\in\{H,L\}$.
We show a stronger property: truthfulness maximizes the expected payoff of $M_{\textrm{DMI}}$ when all other agents play consistent strategies. As a corollary, truthfulness is a Bayes Nash equilibrium even over the larger space of strategies that allow one to choose all $T$ reports simultaneously after looking at all $T$ signals.
Theorem 9.
In the binary signal model, suppose all agents $j\neq i$ play consistent strategies. Then $\sigma_{\text{true}}$ maximizes agent $i$’s expected payment under $M_{\textrm{DMI}}$.
Corollary 6.
The truthful strategy $\sigma_{\text{true}}$ is a Bayes Nash equilibrium of the DMI mechanism.
The key observation needed for Theorem 9 is the following characterization of the possible expected values of each determinant term in the definition of $M_{\textrm{DMI}}$.
Lemma 1.
Suppose agent $j$ plays a consistent strategy. Let $R^{\prime}_{t}$ be the random variable on $\{H,L\}$ resulting agent $j$’s strategy, i.e., $R_{t}^{\prime}\sim\hat{\sigma}(S_{t}^{\prime})$ for each task $t$. Let $p(s_{t})=\Pr[R_{t}^{\prime}=H\mid S_{t}=s_{t}]$, which by assumption is independent of $t$. Then for agent $i$,
$\mathop{\mathbb{E}}[\det M_{12}\mid S_{1},S_{2}]\;=\;\begin{cases}0,&\text{if }r_{1}=r_{2}\text{ or }s_{1}=s_{2},\\ p(H)-p(L),&\text{if }(r_{1},r_{2})=(s_{1},s_{2}),\\ p(L)-p(H),&\text{if }(r_{1},r_{2})=(s_{2},s_{1}).\end{cases}$
where as before
[…]
Proof.
If $r_{1}=r_{2}$, then $M_{12}$ has rank one and $\det M_{12}=0$ deterministically. Suppose next $(r_{1},r_{2})=(H,L)$. We have
$\det M_{12}\;=\;\begin{cases}+1,&\text{if }(R_{1}^{\prime},R_{2}^{\prime})=(H,L),\\ -1,&\text{if }(R_{1}^{\prime},R_{2}^{\prime})=(L,H),\\ 0,&\text{otherwise}.\end{cases}$
As $R_{1}^{\prime},R_{2}^{\prime}$ are independent, we have
$\Pr\bigl{[}(R_{1}^{\prime},R_{2}^{\prime})=(H,L)\mid S_{1},S_{2}\bigr{]}$ $=p(s_{1})\,\bigl{(}1-p(s_{2})\bigr{)}$
$\Pr\bigl{[}(R_{1}^{\prime},R_{2}^{\prime})=(L,H)\mid S_{1},S_{2}\bigr{]}$ $=\bigl{(}1-p(s_{1})\bigr{)}\,p(s_{2})~{}.$
We thus have
$\mathop{\mathbb{E}}[\det M_{12}\mid S_{1},S_{2}]$ $=p(s_{1})\,\bigl{(}1-p(s_{2})\bigr{)}-\bigl{(}1-p(s_{1})\bigr{)}\,p(s_{2})$
$=p(s_{1})-p(s_{2})~{}.$
The same argument gives $\mathop{\mathbb{E}}[\det M_{12}\mid S_{1},S_{2}]=p(s_{2})-p(s_{1})$ when $(r_{1},r_{2})=(L,H)$.
We have now established
$\mathop{\mathbb{E}}[\det M_{12}\mid S_{1},S_{2}]\;=\;\begin{cases}0,&\text{if }r_{1}=r_{2},\\ p(s_{1})-p(s_{2}),&\text{if }(r_{1},r_{2})=(H,L),\\ p(s_{2})-p(s_{1}),&\text{if }(r_{1},r_{2})=(L,H).\end{cases}$
To finish the proof, consider the cases $s_{1}=s_{2}$, $(s_{1},s_{2})=(H,L)$, and $(s_{1},s_{2})=(L,H)$. From the above, we obtain 0 in the first case. In the second, we have $p(H)-p(L)$ if $(r_{1},r_{2})=(s_{1},s_{2})$ and its negation when $(r_{1},r_{2})=(s_{2},s_{1})$. In the third, we again have $p(H)-p(L)$ if $(r_{1},r_{2})=(s_{1},s_{2})$ and its negation when $(r_{1},r_{2})=(s_{2},s_{1})$.
Clearly Lemma 1 also applies to $M_{34}$. And since $\det M_{12}$ and $\det M_{34}$ are independent when conditioned on $S_{1..T}$, the proof of Theorem 9 follows.
Proof of Theorem 9.
Let $s_{1..T}$ be the signal realization for agent $i$, and consider any report $r_{1..T}$. We have
$\mathop{\mathbb{E}}[M_{\textrm{DMI}}(r_{1..T},R^{\prime}_{1..T})\mid S_{1..T}]$ $=\mathop{\mathbb{E}}[\det M_{12}\mid S_{1},S_{2}]\mathop{\mathbb{E}}[\det M_{34}\mid S_{3},S_{4}]~{}.$
$=\mathop{\mathbb{E}}[\det M_{12}\mid S_{1},S_{2}]\mathop{\mathbb{E}}[\det M_{34}\mid S_{3},S_{4}]~{}.$
Applying Lemma 1 to both $M_{12}$ and $M_{34}$, observe that if $s_{1}=s_{2}$ or $s_{3}=s_{4}$ we have $\mathop{\mathbb{E}}[M_{\textrm{DMI}}(r_{1..T},R^{\prime}_{1..T})\mid S_{1..T}]=0$ regardless of $r$. In particular, $\sigma_{\text{true}}$ maximizes the expected payoff.
Now suppose $s_{1}\neq s_{2}$, $s_{3}\neq s_{4}$. Again we have a zero expectation if $r_{1}=r_{2}$ or $r_{3}=r_{4}$. Otherwise, for each matrix, we have two cases: $r$ matches $s$ or not. In particular,
$\mathop{\mathbb{E}}[M_{\textrm{DMI}}(r_{1..T},R^{\prime}_{1..T})\mid S_{1..T}]=\begin{cases}(p(H)-p(L))^{2}&(r_{1},r_{2})=(s_{1},s_{2}),(r_{3},r_{4})=(s_{3},s_{4})\\ -(p(H)-p(L))^{2}&(r_{1},r_{2})\neq(s_{1},s_{2}),(r_{3},r_{4})=(s_{3},s_{4})\\ -(p(H)-p(L))^{2}&(r_{1},r_{2})=(s_{1},s_{2}),(r_{3},r_{4})\neq(s_{3},s_{4})\\ (p(H)-p(L))^{2}&(r_{1},r_{2})=(s_{1},s_{2}),(r_{3},r_{4})=(s_{3},s_{4})\end{cases}$
In summary, the highest expected payoff in this case is $(p(H)-p(L))^{2}$, which is achieved by $\sigma_{\text{true}}$ and the permutation strategy which swaps $H$ and $L$. ∎
Appendix D Omitted Proofs for RBTS
D.1 Gaussian Model
We first prove the form of the functions $\bar{P}(\tau;x)$ and $Q(x)$ for the Gaussian model.
Proposition 11.
In the Gaussian model under the RBTS mechanism,
$\bar{P}(\tau;x)=\Phi\left(\frac{\tau-\rho^{2}x}{\sqrt{b^{2}(1+\rho^{2})(1+\rho)}}\right),$ (20)
and thus,
$Q(x)=\Phi\left(\frac{(1-\rho^{2})\tau}{b\sqrt{(1+\rho)(1+\rho^{2})}}\right).$ (21)
Proof.
Note that for a fixed threshold $\tau$,
$\bar{P}(\tau;x)=\mathop{\mathbb{E}}_{x^{\prime}\sim\Pr[\cdot\mid x]}\Pr[X^{\prime\prime}\leq\tau\mid X^{\prime}=x^{\prime}]=\mathop{\mathbb{E}}_{X^{\prime}\sim N(\rho x,b^{2}(1+\rho))}\Phi\left(\frac{\tau-\rho x^{\prime}}{b\sqrt{1+\rho}}\right),$
and since $X^{\prime}=\rho x+b\sqrt{1+\rho}X$ for $X\sim N(0,1),$
$\bar{P}(\tau;x)$ $=\mathop{\mathbb{E}}_{X}\Phi\left(\frac{\tau-\rho(\rho x+b\sqrt{1+\rho}X)}{b\sqrt{1+\rho}}\right)$
$=\Pr\left[Y\leq\frac{\tau-\rho(\rho x+b\sqrt{1+\rho}X)}{b\sqrt{1+\rho}}\right]$
$=\Pr\left[Y+\rho X\leq\frac{\tau-\rho^{2}x}{b\sqrt{1+\rho}}\right],$
where $Y\sim N(0,1)$. It follows that $Y+\rho X\sim N(0,1+\rho^{2})$, so
$\bar{P}(\tau;x)=\Phi\left(\frac{\tau-\rho^{2}x}{b\sqrt{(1+\rho^{2})(1+\rho)}}\right).$
The definition of $Q(x)=\bar{P}(x;x)$ follows. ∎
Now we have the tools to prove the characterization of equilibria in the Gaussian model.
Proof of Corollary 5.
[…]
$\left(\frac{c_{Q}}{c_{G}}\right)^{2}=\frac{(1-\rho^{2})^{2}}{(1-\rho)^{2}(1+\rho^{2})}=\frac{(1+\rho)^{2}}{1+\rho^{2}}>1.$
Thus $\tau=0$ is the unique point at which $G(\tau)=Q(\tau)$ (see Figure 2 for an example with $a=b=1$), so there are no other potential non-infinite equilibria.
Now we prove $\tau=0$ actually is an equilibrium using the sufficiency condition in Theorem 6. Note that when $\tau=0$, $P(\tau;x)$ and $\bar{P}(\tau;x)$ are both Normal CDFs over $x$ with negative (and distinct) coefficients. One can thus easily check that if $x\leq\tau$, $P(\tau;x)\geq\bar{P}(\tau;x)$. Moreover, if $x>\tau$, $P(\tau;x)\leq\bar{P}(\tau;x)$. It follows by the sufficiency condition in Theorem 6 that $\tau=0$ is an equilibrium. ∎
Next, we include a proof of the form of the best response in RBTS under the Gaussian model.
Proposition 12.
When other agents are playing according to some threshold $\tau$, an agent will best respond with threshold strategy
$\hat{\tau}=\left(\frac{\sqrt{1+\rho^{2}}-1}{\rho(\sqrt{1+\rho^{2}}-\rho)}\right)\tau.$
Proof.
Let $\tau$ be the current threshold strategy that all other agents are following. By equations (18) and (19), an agent who receives signals $x$ will best respond with $H$ if $P(\tau;x)\leq\bar{P}(\tau;x)$, and $L$ otherwise. In the Gaussian model, note that
$P(0;x)$ $=\Phi\left(\frac{-\rho x}{b\sqrt{1+\rho}}\right),$
$\bar{P}(0;x)$ $=\Phi\left(\frac{-\rho^{2}}{b\sqrt{(1+\rho^{2})(1+\rho)}}\right).$
Thus $P(0;x)=\bar{P}(0;x)$ at a unique point which can be verified with simple algebra to be $x=\hat{\tau}$. One can check for the specified value of $\hat{\tau}$, $P(\tau;x)\geq\bar{P}(\tau;x)$ for $x\leq\hat{\tau}$ and $P(\tau;x)\leq\bar{P}(\tau;x)$ for $x\geq\hat{\tau}$. The result follows. ∎
Now we formally compare convergence of the DG and RBTS mechanisms under the Gaussian model.
Proposition 13.
In the Gaussian model, for any initial threshold $\tau(0)\neq 0$, convergence to the stable equilibrium $\tau=0$ is strictly slower in the RBTS mechanism than in the DG mechanism.
Proof.
By Proposition 12, in the RBTS mechanism a best response $\hat{\tau}$ to all other agents playing according to threshold $\tau$ is $\hat{\tau}=m_{\textrm{RBTS}}(\rho)\tau$ for
$m_{\textrm{RBTS}}(\rho)=\frac{\sqrt{1+\rho^{2}}-1}{\rho(\sqrt{1+\rho^{2}}-\rho)}.$
Meanwhile, the best response in DG to a fixed threshold $\tau$ satisfies $P(\tau;\hat{\tau})=F(\tau)$, or
$\Phi\left(\frac{\sqrt{\rho}\tau}{a}\right)=\Phi\left(\frac{\tau-\rho\hat{\tau}}{b\sqrt{1+\rho}}\right).$
The solution to this equality is easily reached by setting the arguments to $\Phi$ equal to each other. We end up with $\hat{\tau}=m_{\textrm{DG}}(\rho)\tau$ for
$m_{\textrm{DG}}(\rho)=\frac{a-b\sqrt{p(1+p)}}{a\rho}.$
Now, one can check that $\frac{m_{\textrm{DG}}(\rho)}{m_{\textrm{RBTS}}(\rho)}>1$ for all $a,b>0$. It immediately follows that the best response $\hat{\tau}$ converges to 0 strictly faster under the DG mechanism than under the RBTS mechanism. ∎
Appendix E Omitted Experiments
We include bifurcation figures for the four-component Gaussian mixture case within the DG and OA mechanisms, as well as bifurcation figures for RBTS. Again, we vary the precision of each component in the Gaussian mixture and study how the stability and number of equilibria change across mechanisms.
(a) Bifurcation graphs for OA (left) and DG (right) in the setting of a four-component Gaussian mixture with means ($-4,-2,2,4$). Following a pattern, we have seven equilibria in OA and five in DG for large enough precision.
(a) 3 component mixture with means (-2, 0, 2)
(b) 4 component mixture with means (-3, -1, 1, 3)
(c) Bifurcation graphs for RBTS in the setting of a three and four-component Gaussian mixture. Note we observe the same stability and number of equilibria as in DG, except that bifurcation occurs sooner.